Circular No. 7318

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999em IN NGC 1637 D. W. Fox and W. H. G. Lewin, Massachusetts Institute of Technology,

report, on behalf of a large collaboration: “SN 1999em has been imaged twice with the Chandra X-ray Observatory (ACIS-S). On Nov. 1–2, 112 photons were detected during a 32-ks observation; preliminary spectral fits indicate a soft source spectrum (photon index ∼ 1.8) with a 0.1–8.0-keV flux of 2 × 10^{−14} erg cm^{−2} s^{−1}, corresponding to an unabsorbed source luminosity of 2 × 10^{38} erg/s at 7.8 Mpc (Sohn and Davidge 1998, A.J. 115, 130). On Nov. 11–12, 60 photons were detected during a 32-ks observation; assuming a (relatively) constant spectrum, this corresponds to a source luminosity of 1 × 10^{38} erg/s.”

C. K. Lacey, Naval Research Laboratory; R. A. Sramek, National Radio Astronomy Observatory (NRAO); S. D. Van Dyk, Infrared Processing and Analysis Center, California Institute of Technology; and K. W. Weiler, Naval Research Laboratory, report: “Radio observations with the NRAO Very Large Array on Nov. 1, 3, 4, 5, 9, 11, 13 and 16 yield nondetections of the type-IIP supernova 1999em. The best upper limits (3σ) from all of the radio observations for SN 1999em are: < 0.86 mJy at 43.3 GHz (7 mm), < 0.77 mJy at 22.5 GHz (1.3 cm), < 0.83 mJy at 15.0 GHz (2 cm) and < 0.26 mJy at 8.4 GHz (3.6 cm). These correspond to upper limits on the radio spectral luminosity of < 6.2, < 5.6, < 6.0 and < 1.9 × 10^{25} erg s^{−1} Hz^{−1}, respectively, for an assumed distance of 7.8 Mpc (Sohn and Davidge, loc. cit.). The full set of measured upper limits is available on

http://rsd-www.nrl.navy.mil/7214/weiler/sne-home.html. Whether

our upper limits imply such a low CSM density that the radio emission is undetectable, or such a high CSM density that it is optically thick to radio emission at this early epoch, remains to be determined through further monitoring.”

Lewin adds: “The combination of x-ray detection and radio nondetection is unusual. The only other known example is the type-IIP SN 1994W in NGC 4041, which has never been detected in the radio (3σ upper limit < 1.1 × 10^{26} erg s^{−1} Hz^{−1} at 8.4 GHz on 1998 Feb. 10) but has a reported ROSAT HRI detection of ∼ 8×10^{39} erg/s during 1997 Oct. 21–23 (Schlegel 1999, Ap. J. in press; astro-ph/9910425). This may imply a significant difference in the physical properties of the presupernova systems between the more readily radio and x-ray detectable type-IIL supernovae, such as 1979C and 1980K, and the type-IIP progenitor systems. We encourage acquisition of additional optical spectra of SN 1999em while its x-ray emission is still strong.”

1999 November 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Brian G. Marsden